# SMP: Reusable Score-Matching Motion Priors for Physics-Based Character Control

### v2

@@ 4. Overview
In this work, we introduce the Score-Matching Motion Prior (SMP), a reusable, modular imitation objective derived from a pretrained diffusion model via score distillation sampling (SDS). The pretrained diffusion model is used to estimate the gradient of the log-likelihood (i.e., the score) of the reference distribution, which evaluates the similarity between an agent’s motions and motions in a reference dataset. The robustness of score-based generative modeling enables a pre-trained diffusion model to provide reliable guidance even for motions that are different from those in the original motion dataset. SMP can be combined with task-specific objectives to train control policies that can accomplish diverse tasks using natural life-like behaviors.
Figure 2 illustrates an overview of our framework. Unlike prior approaches that use diffusion models as planners (serifi2024robot; tevet2024closd; ren2023insactor), SMP repurposes a pretrained diffusion model as a general reward model for evaluating motion naturalness to guide training of a control policy. The task-agnostic diffusion model $f$ is trained solely on reference motion data, independently of any control policy. Once trained, it is frozen and utilized as a reward function via score distillation sampling (SDS). When training a policy, the SDS objective encourages the policy to minimize the discrepancy between the noise $\bm{\epsilon}$ added to the simulated motion, and the noise $\bm{\hat{\epsilon}}$ estimated by the pretrained diffusion model. The SDS error is minimized when the agent’s motions closely aligns with the reference distribution. Similar to other distribution-matching objectives, SMP does not require the simulated character to exactly replicate specific reference motions. Instead, it encourages behaviors that capture the general characteristics of the reference data, enabling smooth transitions and adaptation to tasks that may require skills not explicitly present in the dataset. Furthermore, SMP can be trained on large and diverse datasets by employing a conditional diffusion model. This enables control policies to acquire different stylistic behaviors by conditioning the pretrained diffusion model on style labels, without the need to retain the original motion dataset.
@@ 5. Score-Matching Motion Priors
which is maximized when the simulated motion aligns with the reference distribution. Following standard RL reward design (peng2018deepmimic), we apply an exponential transformation to the SDS loss to normalize the reward between $[0,1]$. While this deviates from the original SDS formulation, we find that this normalization produces empirical improvements when training RL policies.
Previous SDS-based methods have primarily shown promise in the 3D generation, but often produce low-quality, “blurry” results (liang2024luciddreamer). Earlier attempts to use SDS as an objective for reinforcement learning have also struggled to match the performance of adversarial methods for training control policies (luo2024textaware; wu2025diffusing). In the following sections, we introduce several key design choices that improve stability for policy training using SMP, which enable policies to learn high-quality naturalistic human behaviors.
Figure 4. An example of the SDS loss across diffusion noise levels, averaged over 1024 samples. The $y$-axis shows $\log_{10}$ values for better visualization. Larger diffusion timesteps correspond to higher noise levels applied to the sample. The orange region highlights the discrepancy between agent’s motions and training data. SDS loss varies across different diffusion noise levels. At high noise levels, the SDS loss is always minimized, providing limited information about differences between the agent’s motion and the reference distribution, whereas lower noise levels generally yield larger losses. When the diffusion model’s prediction is sufficiently reliable, i.e., on less out-of-distribution (OOD) data, SDS loss computed at lower noise levels can better capture such discrepancies. In our experiments, we compute SDS loss at diffusion timesteps $[22,15,8]$.
@@ 5.1. Ensemble Score-Matching
The SDS objective can be highly sensitive to the choice of diffusion timestep $i$, as different noise scales provide disparate forms of guidance (lin2023magic3d). At higher noise levels, diffused samples are heavily corrupted and dominated by Gaussian noise, which matches the training settings of the diffusion model and leads to more reliable score predictions due to less susceptibility to out-of-distribution samples. Therefore, evaluating the objective at higher diffusion timesteps $i$ can be more instructive when the agent’s motions deviate substantially from the reference data distribution. However, excessive noise also removes much of the information present in the agent behaviors, which may prevent the objective from correcting more minute errors in the agent’s motions. In such cases, the guidance signal may encourage the policy to produce generic or averaged behaviors, as the model no longer retains sufficient information to steer toward more realistic and detailed motions. Therefore, when the difference between the agent’s motions and motions in the dataset is relatively small, diffusing the sample to lower noise levels can better preserve information of the agent’s motions and apply finer-grained corrections.
In prior SDS-based methods in vision domains (poole2022dreamfusion; threestudio2023), it is common to sample the diffusion noise level uniformly $i\sim\mathcal{U}(1,N)$. However, in reinforcement learning, this stochasticity can introduce undesirable variance into the reward signal, leading to less reliable value and advantage estimation. For example, samples diffused with high noise levels are nearly pure Gaussian noise, allowing the diffusion model to predict the noise easily, resulting in small loss values. Whereas samples with lower noise levels generally produce larger denoising errors, as illustrated in Figure 4. This noise-dependent variation in the SDS loss prevents values from randomly sampled timesteps from providing a consistent indicator of the similarity between the agent’s motion and the reference data distribution.
To mitigate this variance, rather than randomly sampling a single timestep per update, we compute a more consistently behaved SDS objective by ensembling multiple SDS evaluations over a fixed set of diffusion timesteps $i\in{\mathbb{K}}$. The resulting SMP reward is:
@@
where $\hat{\bm{\epsilon}}_{i}=f\left(\sqrt{\bar{\alpha}_{i}}\,\tilde{{\mathbf{x}}}^{0}+\sqrt{1-\bar{\alpha}_{i}}\,\bm{\epsilon}_{i}\right)$, $\tilde{{\mathbf{x}}}$ denotes the simulated character’s motion, and $\mathbb{K}$} is a predefined set of diffusion timesteps. In all experiments, we use ${\mathbb{K}}=\{0.44N,0.30N,0.16N\}$, where $N$ is the total number of diffusion steps.
@@
5.2. Generative State Initialization
Reference state initialization (RSI), where the simulated character is initialized to states randomly sampled from a reference motion dataset, has been shown to be a vital technique for improving exploration in motion imitation methods (peng2018deepmimic). However, RSI requires access to the motion dataset to sample initial states from during policy training. To remove this reliance on the original dataset, SMP can instead leverage the generative prior to generate initial states when training the policy. When using SMP to train a policy, initial states are generated by sampling from the diffusion model used as the pretrained motion prior. SMP therefore serves dual purposes in our framework, as both a reward function and an initial state distribution when training new policies. This generative state initialization (GSI) method alleviates the need to retain the original motion dataset once the SMP has been trained by leveraging the generative capabilities of diffusion models to produce diverse, high-quality initial states.
6. Model Representation
@@ 6.1. Motion Representation
Designing an appropriate motion representation is critical both for training the diffusion model to capture the dataset distribution, and for repurposing it as a motion prior for policy training. The representation should contain sufficient information to reconstruct motion while remaining easy to extract from both kinematic motion clips and simulator state observations. Following the design of prior works in motion diffusion and AMP (tevet2023human; zhang2022motiondiffuse; peng2021amp), our motion features include:
@@
The character’s local coordinate frame is defined with the origin at the root (i.e., pelvis), the $x$-axis aligned with the root link’s facing direction, and the $y$-axis aligned with the global up vector. Joint rotations are represented using a 6D representation for spherical joints (zhou2019continuity).
@@ 6.2. Diffusion Model Implementation
The motion diffusion model is implemented as a transformer encoder, using adaptive normalization to inject noise-level conditions (and style conditions, when available). Unless otherwise specified, we use a window of $n=10$ frames. We find that a carefully designed two-layer transformer encoder with only 3M parameters is sufficient to capture the distribution of large-scale motion datasets, such as 100STYLE (mason2022real) with over 20 hours of stylized motions. The number of diffusion timesteps is set to $N=50$, and the model is trained to predict $\bm{\epsilon}$. Following standard practice for training diffusion models, we apply exponential moving average (EMA) on the model parameters during training (karras2024analyzing; dhariwal2021diffusion). The model is trained for 400k–800k iterations, depending on the dataset, with convergence typically achieved within five hours on a single RTX 4090 GPU.
6.3. Policy Implementation
Following prior work, the control policy $\pi$ is modeled using a multi-layer perceptron (MLP) that maps an input state ${\mathbf{s}}_{t}$ and goal $\mathbf{g}$} to a Gaussian distribution over the action space ${\mathbf{a}}_{t}\in\mathcal{A}$ (peng2018deepmimic). The value function $V({\mathbf{s}}_{t},{\mathbf{g}})$ is modeled by a separate MLP network with a similar architecture to the policy.
@@ 6.3. Policy Implementation
The state ${\mathbf{s}}_{t}$ is represented using features similar to those in peng2021amp, such as body link positions, link rotations encoded in a 6D representation, and linear and angular velocities of each link. All features are calculated in the character’s local coordinate frame. Since our policies are not trained to track specific reference motions, no phase variable or target reference state information is provided in the state. The action ${\mathbf{a}}_{t}$ specifies target joint positions for PD controllers at each joint. For spherical joints, each target rotation is parameterized using a 3D exponential map $q\in\mathbb{R}^{3}$ (grassia1998practical).
@@
The policy is trained using proximal policy optimization (PPO) (schulman2017proximal). At each timestep $t$, the agent queries the motion prior (see Algorithm 1) to obtain a prior reward $r_{t}^{\mathrm{smp}}$, and may also receive a task reward $r_{t}^{g}$ from the environment. These are combined linearly to form the composite reward at that timestep
@@
following peng2021amp. After collecting a batch of trajectories, mini-batches are sampled from the buffer to update both the policy and the value function. The policy is updated with PPO using advantages estimated via $\mathrm{GAE}(\lambda)$ (schulman2015high), while the value function is trained with targets computed via $\mathrm{TD}(\lambda)$ (sutton1998reinforcement). The pretrained diffusion model is kept fixed during the policy training process, and does not need to be updated.
(a) Target Location
(b) Steering
(c) Dodgeball
@@
Figure 5. Score-matching motion priors can be trained on datasets of varying sizes, independently of any task or control policy. Once trained, an SMP provides a motion imitation objective $r^{\mathrm{smp}}$, which can be composed with task rewards $r^{\mathrm{g}}$ to train multiple policies that complete a diverse array of tasks while exhibiting natural, life-like behaviors.
@@ 7. Tasks
We evaluate the effectiveness of SMP across six motion control tasks, showcasing its ability to train policies that can perform various control tasks using diverse motion styles. Below, we summarize each task and the corresponding goal observation $\mathbf{g}$}, which provide the agent with task-relevant information from the environment. More comprehensive details and task reward functions are provided in Appendix.
@@
In this task, a ball is launched toward the character from a random position up to $10\mathrm{m}$ away, with a launch speed sampled from $20,25]\mathrm{m/s}$. This gives the agent less than $0.5\mathrm{s}$ to react and dodge. If the character is hit, the episode terminates early and the agent receives zero reward for all remaining timesteps as a penalty. The agent receives ball information through the goal vector ${\mathbf{g}}_{t}=\left({\mathbf{p}}_{t}^{\mathrm{ball}},\dot{{\mathbf{p}}}_{t}^{\mathrm{ball}}\right).$
@@
In addition to training priors for human motion, SMP can also be used to train priors that jointly model the interactions between humans and objects. First, we train an SMP on a dataset of human-object carrying motions. This prior is then used to train policies for an object carrying task, where the character is required to carry a box from the ground to a randomly placed target location. The goal ${\mathbf{g}}_{t}={{\mathbf{p}}^{\mathrm{box}}_{t}}^{*}$ records the target box position. The state ${\mathbf{s}}_{t}$ is augmented with additional features that describe the state of the box, including the position ${\mathbf{p}}^{\mathrm{box}}_{t}$ and the orientation ${\mathbf{q}}^{\mathrm{box}}_{t}$. All box information is represented in the character’s local coordinate.
@@
In this task, the policy trained to recover from arbitrary fallen states and maintain a minimum root height of $0.8m$.
8. Results
To evaluate the effectiveness of SMP, we apply our framework to train policies for on a suite of challenging motion control tasks. In Section 8.1, we demonstrate themodularity of SMP by using a pre-trained prior to train new control policies without requiring access to the original motion dataset. By training a single style-conditioned diffusion model on the 20-hour 100STYLE dataset, the prior can then be used to train policies to perform tasks in a variety of different styles, where each style-specific prior reward is obtained simply by conditioning the pretrained model on the desired style label. New styles can also be synthesized by composing the pretrained diffusion model when conditioned on different styles. In Section 8.2, we show that a single motion prior can be reused to train policies across diverse tasks, such as steering, target location, and dodgeball. In addition to controlling the behavioral style of the character, in Section 8.3 we show that SMP can also model human-object interaction priors. By training the diffusion model on human-object interaction data, SMP learns to jointly model both object and character motions, allowing agents to use human-like behavior to perform object interaction tasks. SMP also enables agents to learn robust and natural recovery behaviors, as demonstrated in Section 8.4. Similar to other distribution-matching objectives, the policies trained with SMP can synthesize new skills that are not explicitly present in the reference dataset. In Section 8.5, when trained on a 3-second walk–jog–run dataset, SMP automatically leads to the emergence of different locomotion gaits that enable a character to closely follow a wide spectrum of target speeds, as well as natural transitions between the various gaits. Finally, to benchmark SMP’s effectiveness for motion imitation, we evaluate SMP on single-clip imitation tasks (Section 8.6). SMP is able to closely reproduce a diverse set of dynamic and acrobatic skills, achieving comparable motion quality to state-of-the-art adversarial imitation learning methods.

### v3

@@ 4. Overview
In this work, we introduce the Score-Matching Motion Prior (SMP), a reusable, modular imitation objective derived from a pretrained diffusion model via score distillation sampling (SDS). The pretrained diffusion model is used to estimate the gradient of the log-likelihood (i.e., the score) of the reference distribution, which evaluates the similarity between an agent’s motions and motions in a reference dataset. The robustness of score-based generative modeling enables a pre-trained diffusion model to provide reliable guidance even for motions that are different from those in the original motion dataset. The general style imitation objective provided by SMP can be combined with task-specific objectives to train control policies that can accomplish diverse tasks using natural life-like behaviors.
Figure 2 illustrates an overview of our framework. Unlike prior approaches that use diffusion models as planners (Serifi et al., 2024; Tevet et al., 2024; Ren et al., 2023), SMP repurposes a pretrained diffusion model as a general reward model for evaluating motion naturalness to guide training of a control policy. The task-agnostic diffusion model $f$ is trained solely on reference motion data, independently of any control policy or task. Once trained, it is frozen and utilized as a reward function via score distillation sampling (SDS). When training a policy, the SDS objective encourages the policy to minimize the discrepancy between the noise $\bm{\epsilon}$ added to the simulated motion, and the noise $\bm{\hat{\epsilon}}$ estimated by the pretrained diffusion model. The SDS error is minimized when the agent’s motions closely aligns with the reference distribution. Similar to other distribution-matching objectives, SMP does not require the simulated character to exactly replicate specific reference motions. Instead, it encourages behaviors that capture the general characteristics of the reference data, enabling smooth transitions and adaptation to tasks that may require skills not explicitly present in the dataset. Furthermore, SMP can be trained on large and diverse datasets by employing a conditional diffusion model. This enables policies to acquire different stylistic behaviors by conditioning the pretrained diffusion model on style labels, without the need to retain the original motion dataset.
@@ 5. Score-Matching Motion Priors
which is maximized when the simulated motion aligns with the reference distribution. Following standard RL reward design (Peng et al., 2018a), we apply an exponential transformation to the SDS loss to normalize the reward between $[0,1]$. While this deviates from the original SDS formulation, we find that this normalization produces empirical improvements when training RL policies.
Previous SDS-based methods have primarily shown promise in the 3D generation, but often produce low-quality, “blurry” results (Liang et al., 2024). Earlier efforts to use SDS as a reinforcement learning objective have also struggled to match the performance of adversarial methods for training control policies (Luo et al., 2024; Wu et al., 2025a). In the following sections, we introduce several key design decisions that improve stability for policy training using SMP, enabling policies to learn high-quality naturalistic behaviors.
Figure 4. An example of the SDS loss across diffusion noise levels. The $y$-axis shows $\log_{10}$ values for visualization. Larger diffusion timesteps correspond to higher noise levels applied to the sample. Statistics are averaged across 1024 samples for each diffusion timestep. The orange region highlights the difference between the SDS scores from the agent’s motions and the motion data. The scale of the SDS loss can vary significantly across different diffusion noise levels.
@@ 5.1. Ensemble Score-Matching
The SDS objective can be highly sensitive to the choice of diffusion timestep $i$, as different noise scales provide disparate forms of guidance (Lin et al., 2023). At higher noise levels, diffused samples are heavily corrupted and dominated by Gaussian noise, which matches the training settings of the diffusion model and leads to more reliable score predictions due to less susceptibility to out-of-distribution samples. Therefore, evaluating the objective at higher diffusion timesteps $i$ can be more instructive when the agent’s motions deviate substantially from the reference data distribution. However, excessive noise also removes much of the information present in the agent behaviors, which may prevent the objective from correcting more minute errors in the agent’s motions. In such cases, the guidance signal may encourage the policy to produce generic or averaged behaviors, as the model no longer retains sufficient information to steer toward more realistic and detailed motions. Therefore, when the difference between the agent’s motions and motions in the dataset is relatively small, diffusing the sample to lower noise levels can better preserve information of the agent’s motions and apply finer-grained corrections.
In prior SDS-based methods in vision domains (Poole et al., 2022; Guo et al., 2023), it is common to sample the diffusion noise level uniformly $i\sim\mathcal{U}(1,N)$. However, in reinforcement learning, this stochasticity can introduce undesirable variance into the reward signal, leading to less reliable value and advantage estimation. For example, samples diffused with high noise levels are nearly pure Gaussian noise, allowing the diffusion model to easily predict the applied noise, leading to small SDS errors and limited information about differences between the agent’s motions and the reference distribution, as illustrated in Figure 4. Whereas samples with lower noise levels generally produce larger SDS errors. Due to this noise-dependent variation in the SDS loss, computing the SMP reward with randomly sampled timesteps can result in a highly stochastic and unreliable measure of the similarity between the agent’s motions and the reference data distribution.
To reduce the variance of the SMP reward function, instead of randomly sampling a single timestep per reward calculation, we propose ensemble score-matching (ESM), which constructs a lower variance reward function by aggregating multiple SDS evaluations over a fixed set of diffusion timesteps $i\in{\mathbb{K}}$. The resulting SMP reward is given by:
@@
where $\hat{\bm{\epsilon}}_{i}=f\left(\sqrt{\bar{\alpha}_{i}}\,\tilde{{\mathbf{x}}}^{0}+\sqrt{1-\bar{\alpha}_{i}}\,\bm{\epsilon}_{i}\right)$, $\tilde{{\mathbf{x}}}$ denotes the simulated character’s motion, and $\mathbb{K}$} is a predefined set of diffusion timesteps. When the diffusion model’s predictions are sufficiently reliable, i.e., for samples that are less out-of-distribution (OOD), the SDS loss computed at lower noise levels responds more sensitively to discrepancies between the agent’s motion and the reference training distribution. In our experiments, we found that SDS error computed at low diffusion timesteps is highly sensitive to jittery artifacts, whereas SDS error computed at medium timesteps is more robust and provides better guidance for correcting behavioral artifacts. Moreover, evaluating a large number of diffusion steps incurs significant computational cost, and adjacent timesteps tend to provide redundant guidance at similar granularity levels. We therefore found that evaluating the SMP objective using a small set of representative diffusion timesteps is sufficient for effective performance. In all experiments, we compute the SDS loss at ${\mathbb{K}}=\{22,15,8\}$, which provides an effective balance across different noise scales.
@@
5.2. Generative State Initialization
Reference state initialization (RSI), where the simulated character is initialized to states randomly sampled from a reference motion dataset, has been shown to be a vital technique for improving exploration in motion imitation methods (Peng et al., 2018a). However, RSI requires access to the motion dataset to sample initial states from during policy training. To remove this reliance on the original dataset, SMP can instead leverage the generative prior to generate initial states when training the policy. When using SMP to train a policy, initial states are generated by sampling from the diffusion model used as the pretrained motion prior. SMP therefore serves dual purposes in our framework, as both a reward function and an initial state distribution when training new policies. This generative state initialization (GSI) method alleviates the need to retain the original motion dataset once the SMP has been trained by leveraging the generative capabilities of diffusion models to produce diverse, high-quality initial states.
6. Model Representation
@@ 6.1. Motion Representation
Designing an appropriate motion representation is critical both for training the diffusion model to capture the dataset distribution, and for repurposing it as a motion prior for policy training. The representation should contain sufficient information to reconstruct motion while remaining easy to extract from both kinematic motion clips and simulator state observations. Following the design of prior works in motion diffusion and AMP (Tevet et al., 2023; Zhang et al., 2022; Peng et al., 2021), our motion features include:
@@
The character’s local coordinate frame is defined with the origin at the root (i.e., pelvis), the $x$-axis aligned with the root link’s facing direction, and the $y$-axis aligned with the global up vector. Joint rotations are represented using a 6D representation for spherical joints (Zhou et al., 2019).
@@ 6.2. Diffusion Model Implementation
The motion diffusion model is implemented as a transformer encoder, using adaptive normalization to inject noise-level conditions (and style conditions, when available). Unless otherwise specified, we use a window of $H=10$ frames. We find that a simple two-layer transformer encoder with only 3M parameters is sufficient to capture the distribution of large-scale motion datasets, such as 100STYLE (Mason et al., 2022) with over 20 hours of stylized motions. The number of diffusion timesteps is set to $N=50$, and the model is trained to predict $\bm{\epsilon}$. Following standard practice for training diffusion models, we apply exponential moving average (EMA) on the model parameters during training (Karras et al., 2024; Dhariwal and Nichol, 2021). The model is trained for 400k–800k iterations, depending on the dataset, with convergence typically achieved within five hours on a single RTX 4090 GPU.
6.3. Model Implementation
Following prior work, the control policy $\pi$ is modeled using a multi-layer perceptron (MLP) that maps an input state ${\mathbf{s}}_{t}$ and goal $\mathbf{g}$} to a Gaussian distribution over the action space ${\mathbf{a}}_{t}\in\mathcal{A}$ (Peng et al., 2018a). The value function $V({\mathbf{s}}_{t},{\mathbf{g}})$ is modeled by a separate MLP network with a similar architecture to the policy.
@@ 6.3. Model Implementation
The state ${\mathbf{s}}_{t}$ is represented using features similar to those in Peng et al. (2021), such as body link positions, link rotations encoded in a 6D representation, and linear and angular velocities of each link. All features are calculated in the character’s local coordinate frame. Since our policies are not trained to track specific reference motions, no phase variable or target reference state information is provided in the state. The action ${\mathbf{a}}_{t}$ specifies target joint positions for PD controllers at each joint. For spherical joints, each target rotation is parameterized using a 3D exponential map $q\in\mathbb{R}^{3}$ (Grassia, 1998).
@@
The policy is trained using proximal policy optimization (PPO) (Schulman et al., 2017). At each timestep $t$, the agent queries the motion prior (see Algorithm 1) to obtain a prior reward $r_{t}^{\mathrm{smp}}$, and may also receive a task reward $r_{t}^{g}$ from the environment. These are combined linearly to form the composite reward at that timestep
@@
following Peng et al. (2021). After collecting a batch of trajectories, mini-batches are sampled from the buffer to update both the policy and the value function. The policy is updated with PPO using advantages estimated via $\mathrm{GAE}(\lambda)$ (Schulman et al., 2015), while the value function is trained with targets computed via $\mathrm{TD}(\lambda)$ (Sutton et al., 1998). The pretrained diffusion model is kept fixed during the policy training process, and does not need to be updated.
Training a policy with SMP on a single motion clip (e.g., spinkick) takes approximately 30 minutes on a single RTX 4090 GPU. Training a HighKnees policy for the target location task takes 11.5 hours for 600M samples, compared to 6.2 hours for AMP. SMP uses a larger prior model that captures 100 styles, whereas AMP uses a smaller style-specific discriminator. Runtime performance is identical to AMP, since only the policy is executed.
(a) Dodgeball
(b) Target Location
(c) Steering
@@
Figure 5. Score-matching motion priors can be trained on a wide range of dataset sizes, independently of any task or control policy. Once trained, an SMP provides a motion imitation objective $r^{\mathrm{smp}}$ that can be composed with task rewards $r^{\mathrm{g}}$ to train policies that perform a diverse array of tasks, all the while exhibiting natural, life-like behaviors.
@@ 7. Tasks
We evaluate the effectiveness of SMP across seven motion control tasks, showcasing its ability to train policies that can perform various control tasks using diverse motion styles. Below, we summarize each task and the corresponding goal observation $\mathbf{g}$}, which provides the agent with task-relevant information from the environment. More comprehensive details, including task reward functions and specific hyperparameter settings are provided in Appendix D.
@@
In this task, a ball is launched toward the character from a random position up to $10\mathrm{m}$ away, with a launch speed sampled from $20,25]\mathrm{m/s}$. This gives the agent less than $0.5\mathrm{s}$ to react and dodge. If the character is hit, the episode terminates early and the agent receives zero reward for all remaining timesteps as a penalty. The agent observes the ball state through goal observations ${\mathbf{g}}_{t}=\left({\mathbf{p}}_{t}^{\mathrm{ball}},\dot{{\mathbf{p}}}_{t}^{\mathrm{ball}}\right)$, where ${\mathbf{p}}_{t}^{\mathrm{ball}}$ and $\dot{{\mathbf{p}}}_{t}^{\mathrm{ball}}$ denote the ball’s position and velocity represented in the character’s local coordinate frame.
@@
In addition to training priors for human motion, SMP can also be used to train priors that jointly model the interactions between the character and objects. First, we train an SMP on a dataset of human-object carrying motions. This prior is then used to train policies for an object carrying task, where the character is required to carry a box from the ground to a randomly placed target location. The goal ${\mathbf{g}}_{t}={{\mathbf{p}}^{\mathrm{box}}_{t}}^{*}$ records the target box position. The state ${\mathbf{s}}_{t}$ is augmented with additional features that describe the state of the box, including the position ${\mathbf{p}}^{\mathrm{box}}_{t}$ and the orientation ${\mathbf{q}}^{\mathrm{box}}_{t}$. All features are represented in the character’s local coordinate frame.
Stair Traversal
To further evaluate SMP’s ability for modeling character-object interactions, we apply SMP to train policies for a stair traversal task. The task requires the character to traverse a short staircase by walking up and then down a series of steps while following a target velocity $\mathbf{v}^{\text{tar}}$. At each timestep, the goal $\mathbf{g}_{t}=\hat{\mathbf{v}}^{\text{tar}}_{t}$ records the target velocity expressed in the character’s local coordinate frame. The state ${\mathbf{s}}_{t}$ is augmented with additional features that encode the staircase’s position and orientation.
@@
In this task, the policy is trained to recover from arbitrary fallen states and maintain a standing pose.
8. Experiments
To evaluate the effectiveness of SMP, we apply our framework to train policies for a suite of challenging motion control tasks. In Section 8.1, we demonstrate the modularity of SMP by using a pre-trained prior to train new control policies without requiring access to the original motion dataset. By training a style-conditioned diffusion model on the 20-hour 100STYLE dataset, we show that a single pre-trained prior can then be used to train policies to perform tasks in a variety of different styles, where each style-specific prior reward is obtained simply by conditioning the pretrained SMP on the desired style label. New styles can also be synthesized by composing the pretrained diffusion model’s predictions from different styles. In Section 8.2, we show that a single motion prior can be reused to train policies across diverse tasks, such as steering, target location, and dodgeball. In addition to guiding the behavioral style of the character, in Section 8.3 we show that SMP can also model human-object interactions. By training the diffusion model on human-object interaction data, SMP learns to jointly model both object and character motions, allowing agents to use human-like behavior to perform object interaction tasks. Beyond task-directed behaviors, SMP also enables agents to learn robust and natural recovery behaviors, as demonstrated in Section 8.4. Similar to prior distribution-matching objectives, the policies trained with SMP can synthesize new skills that are not explicitly present in the reference dataset. In Section 8.5, we show that an SMP trained on a small-scale dataset, containing only 3 seconds of walking, jogging, and running motion clips, can automatically lead to the emergence of different locomotion gaits that enable a character to closely follow a wide range of different target speeds, as well as natural transitions between various gaits. Finally, to benchmark SMP’s effectiveness for motion imitation, we evaluate SMP on single-clip imitation tasks (Section 8.6). SMP is able to closely reproduce a diverse set of dynamic and acrobatic skills, achieving comparable motion quality to state-of-the-art adversarial imitation learning methods.
Figure 6. A pretrained 100-style motion prior can be adapted to synthesize motion priors of new styles. For example, we create a novel “AeroPlane + HighKnees” prior by blending two existing styles in the $\bm{\epsilon}$-space. This crafted prior, which is used for both generative state initialization and the motion prior objective $r^{\mathrm{smp}}$, enables the agent to perform the target location task with a new, agile style, all without requiring the corresponding reference data.